OpenStudy (anonymous):

Which ratio is equivalent to 7 : 9? nine over seven 14 to 18 21 : 3 7 to 27

OpenStudy (anonymous):

Which is the only that makes sense here?

OpenStudy (anonymous):

@lupita1995

OpenStudy (anonymous):

Ratio means to divide...

OpenStudy (fibonaccichick666):

batman... are you implying what I think?

OpenStudy (fibonaccichick666):

pm me if so

OpenStudy (anonymous):

Mhm?

OpenStudy (anonymous):

21:3??

OpenStudy (anonymous):

\[\frac{ 7 }{ 9 } = \frac{ 21 }{ 3 }?\] Really?

OpenStudy (anonymous):

omg its a fudging question

OpenStudy (anonymous):

I know you're looking for just the answer, so Idc.

OpenStudy (fibonaccichick666):

work is usually appreciated on here

OpenStudy (anonymous):

I'm not just looking for the answer ur supposed to help and u get fudging angry

OpenStudy (fibonaccichick666):

otherwise we cannot assist in the learning process

OpenStudy (anonymous):

I will help soon as you put some effort, what is 7/9 = ?

OpenStudy (anonymous):



OpenStudy (fibonaccichick666):

can you type what this means, \(\frac{7}{9}\) , then tell me what \(7\lceil{9}\) and \(9\lceil{7}\)means?

OpenStudy (anonymous):

what??

OpenStudy (anonymous):